# Multiple Imputation of Hierarchical Nonlinear Time Series Data with an Application to School Enrollment Data

### v1

@@ 4.4 Model Implementation
For each replication of each experiment, we created 40 imputations using the MINTS method by running 10 chains of the MCMC algorithm. The bounds of the model for $\mathbf{X}$ were set as $X_{low}=0$ and $X_{up}=100$ and the bounds of the model for $\mathbf{Y}|\mathbf{X}$ were set as $Y_{low}=0$ and $Y_{up}=\text{min}(X_{c,t},60)$. During the estimation phase, the MCMC was run for enough iterations to ensure convergence of all imputation model parameters. The number of iterations differed across experiments, but ranged from 10000 to 25000 iterations per chain. During the imputation phase, an additional 4000 iterations was run for each chain and four iterations from each chain were selected as the imputed values. The iterations selected as the final imputed values were chosen to be 1000 iterations apart to ensure autocorrelation was close to zero following the procedure described in Section 3.3.2.
@@
We compared the MINTS method to six models based on existing methods for multiple imputation. Three of these models are based on the MICE methodology as implemented in the R package “mice” (van Buuren and Groothuis-Oudshoorn, 2011).222mice version 3.14.0 used MICE uses the FCS algorithm, which allows for the specification of separate univariate conditional models for each variable with missing data. The univariate conditional models include all available variables as predictors, for example, the model for $\mathbf{X}$ includes $\mathbf{Y}$, year, and country as predictors. We first considered the default imputation method for continuous data in the mice package, which is predictive mean matching. We refer to this method as the MICE PMM method. Unlike the other methods considered, MICE PMM does not explicitly account for the hierarchical structure of the data but is included as a baseline for comparison.
@@
We evaluated two models within the MICE framework that account for the hierarchical structure of the data using the method of Schafer (1997b) and Schafer and Yucel (2002), referred to as the pan method following its implementation in the R package “pan” (Zhao and Schafer, 2023). The function mice.impute.2l.pan within the mice package uses a Gibbs sampler to estimate the conditional linear mixed effects model with homogeneous within-group variances for each variable with missing data given the other variables. We evaluated a model that includes a country-specific intercept and fixed effects for all covariates in the imputation model, referred to as the pan Fixed Effects method. For example, imputed values for $X_{c,t}|Y_{c,t}$ are modeled with a country-specific intercept, a fixed effect of year, a fixed effect of $Y_{c,t}$, and a homogeneous normally distributed error term. We also considered the pan Random Effects method, which adds random effects of all covariates to the model used in the pan Fixed Effects method.
@@
We evaluated three models using the Amelia II methodology as implemented in the R package “Amelia” (Honaker and King, 2010). [^3] Amelia version 1.80 used For complete data $(\mathbf{X},\mathbf{Y})$, Amelia assumes the joint distribution is multivariate normal. The parameters of the joint distribution are estimated using a combination of bootstrapping and an EM algorithm. Amelia is designed specifically for imputation of hierarchical time series data, which Honaker and King refer to as time series cross-sectional data, through modeling features such as smooth trends over time and allowing for country-specific effects. For our comparisons, we evaluated three implementations of Amelia that were chosen to use the simplest form of the time-series-cross-sectional modeling features in Amelia: a time-series (TS) model, a cross-sectional (CS) model, and a time-series-cross-sectional (TSCS) model. All three of the imputation models are constructed by adding terms to the default multivariate normal joint model for $(\mathbf{X},\mathbf{Y})$. In the Amelia TS method, a linear effect of time is added. In the Amelia CS method, a country-specific intercept term is added. In the Amelia TSCS method, a country-specific intercept term and a country-specific linear effect of time are added.
@@ 4.5 Analysis Model Validation Results
For the linear regression analysis model validation, we evaluated how well each multiple imputation method performs for estimation of $Q=\omega_{1}$, the regression coefficient on $\mathbf{Y}$ in the linear regression of $\mathbf{Z}$ on $\mathbf{Y}$. Table 1 summarizes the results of this validation for the nonlinear simulated data. Overall, MINTS results in the best balance between MAE, coverage, and FMI across the nine experiments. MINTS has the smallest MAE in all experiments except MAR 10%, where MINTS has the second smallest MAE behind MICE PMM. At the 10% and 40% rates, MINTS has reasonably close to nominal coverage for 95% intervals. While MINTS has the closest to nominal coverage out of the methods compared in the experiments at the 80% rate, coverage is below nominal in all three experiments. MINTS also has the smallest FMI at the 10% and 40% rates. As expected, FMI is much larger for all methods at the 80% rate, with FMI $>50\%$ for several methods. This is higher than is typical for the setting of sample survey data that multiple imputation was originally designed for, where FMI is usually $\leq 30\%$ (Rubin, 2003). Given the high FMI, it is thus unsurprising how poorly all multiple imputation methods perform in the experiments at the 80% rate.
None of the existing methods perform consistently well across experiments for estimation of $\omega_{1}$. While MICE PMM can outperform the methods explicitly designed for hierarchical time series data at the lower rates of missingness, the performance of MICE PMM suffers greatly at the highest rate of missingness. The pan and Amelia methods perform similarly to one another at the 10% rate, but all have larger MAE than MICE PMM. Performance of the pan and Amelia methods generally is best under MCAR, with pan Random Effects and Amelia TSCS performing the best of the group in terms of MAE and coverage.
@@
Simulated Method MCAR MAR MNAR Missingness Rate MAE Cvg FMI MAE Cvg FMI MAE Cvg FMI 10% MICE PMM 1.16 100.0 6.4 0.53 100.0 3.4 0.81 100.0 4.4 pan Fixed 1.78 100.0 6.0 6.34 93.3 7.1 3.43 100.0 5.5 pan Random 1.18 100.0 4.1 4.87 99.9 4.9 2.35 100.0 3.5 Amelia TS 1.87 100.0 7.1 5.64 99.8 7.6 3.37 100.0 6.0 Amelia CS 2.30 100.0 8.1 6.83 95.3 9.1 3.97 100.0 6.8 Amelia TSCS 1.28 100.0 4.7 5.13 99.9 5.2 2.45 100.0 3.6 MINTS 0.29 100.0 0.5 0.56 100.0 0.5 0.29 100.0 0.4 40% MICE PMM 13.76 30.5 37.8 4.71 100.0 23.0 10.25 56.7 32.3 pan Fixed 10.35 50.7 24.9 27.75 0.0 22.7 17.86 0.0 22.2 pan Random 4.95 97.5 16.6 20.05 0.0 15.9 9.95 16.9 14.7 Amelia TS 11.77 51.0 23.8 23.35 0.0 21.3 16.87 0.0 21.7 Amelia CS 19.61 1.0 29.0 27.31 0.0 22.2 22.57 0.0 24.3 Amelia TSCS 5.19 96.9 22.3 21.23 0.0 23.8 9.94 21.3 17.4 MINTS 1.21 100.0 3.9 3.02 99.9 4.4 1.49 100.0 2.8 80% MICE PMM 60.62 0.1 60.2 29.59 4.8 56.9 59.98 0.0 54.4 pan Fixed 35.58 2.9 49.0 56.03 0.0 35.8 49.51 0.0 43.6 pan Random 15.30 43.2 54.1 43.08 0.0 51.7 26.93 0.1 52.7 Amelia TS 41.08 4.8 50.4 54.29 0.0 32.9 46.55 0.0 33.1 Amelia CS 99.19 0.0 66.6 79.16 0.0 70.0 77.02 0.0 54.4 Amelia TSCS 22.73 48.2 75.8 57.15 0.0 78.9 35.91 4.9 76.6 MINTS 7.62 81.0 38.7 13.14 47.5 47.0 9.71 62.6 35.6
We note that the existing methods included in the comparisons assume a linear relationship between variables in the imputation model. We also conducted a simulation study for data simulated to follow a linear relationship where we considered two uncongenial and two congenial analysis models. Although our primary focus is on the uncongenial setting, the analysis model validation using congenial analysis models and linear simulated data allows us to compare the performance of MINTS with the existing imputation methods in an “ideal” setting for multiple imputation. We found that the existing imputation methods perform better in the analysis model validation for the linear simulated data compared to the nonlinear simulated data. However, for estimation of uncongenial analysis models we found MINTS still outperforms the existing methods in the linear setting at the 10% and 40% rates of simulated missingness. All imputation methods were found to perform well for estimation of congenial analysis models using linear simulated data, with no method consistently standing out from the others across experiments. Details of the simulation study using linear simulated data can be found in the Appendix.
@@ 5 Application to Enrollment Data
We further evaluated the performance of MINTS by revisiting the secondary school enrollment rate data that was described in Section 2. We conducted two validation exercises using the enrollment data by simulating additional country-years as missing. In the first validation exercise, we evaluated how well MINTS performs for estimation of parameters of interest $Q$ for uncongenial analysis models. This is analogous to the analysis model validation that was conducted for the simulation study.
In the second validation exercise, we evaluated the predictive performance of MINTS for predicting left-out observations of NER. Out-of-sample validation for prediction of individual missing values is less frequently conducted for multiple imputation methods compared to analysis model validation, but has been considered by Gelman et al. (1998); Honaker and King (2010); Nguyen et al. (2017); among others. Although the primary goal of multiple imputation is to create valid estimates of parameters of interest in the presence of missing data rather than recovering the missing values (Rubin, 1996), the prediction of individual missing values can still be of great interest in practice. A multiple imputation method that can perform well for prediction of individual missing values and for creating multiply imputed estimates of quantities of interest is thus of increased utility. We refer to this second validation exercise as “out-of-sample validation.”
Finally, we applied the MINTS method to the full enrollment data set without simulating any missingness and created 40 multiple imputations for the country-years of NER and GER that are missing in the original data set. We briefly present the multiple imputation results and make available the 40 completed data sets.
@@ 5.1 Validation Procedure
For both validation exercises using the enrollment data, we considered eight experimental conditions. Parameters varied in both validation exercises were the rate of simulated missingness (10%, 40%, 80%) and the missing data mechanism (MCAR, MAR, and MNAR), where the same missing data mechanisms used in the simulation study and described in the Appendix were also used for the enrollment data. We did not consider the MNAR 80% experiment for the enrollment data as this resulted in the majority of countries having no observations for NER or GER.
@@
For the analysis model validation, each experiment was replicated $N_{rep}=100$ times following an analogous procedure as was described in Section 4.3 for the simulation study. A major difference in the analysis model validation procedure for the enrollment data is the need to distinguish between the country-years that start out as missing and the country-years that are simulated as missing, where only the country-years that are simulated as missing are used for evaluation purposes. To facilitate this distinction, in each replication of each experiment we separated the enrollment data into “started-as-missing”, “observed”, and “simulated-as-missing” sets. The observed set is the training set for model estimation, while the simulated-as-missing set is the testing set for validation. Imputed values were still created for the started-as-missing set, but we are unable to evaluate the performance of these imputed values as the true value for the country-years that started as missing is unknown.
For the out-of-sample validation exercise, we considered one replication of each experiment. We used the same distinction between the country-years that started as missing and the country-years that were simulated as missing, where only the country-years that were simulated as missing are included in the testing set for validation. The out-of-sample validation compares performance metrics for prediction of missing values for NER averaged over all country-years in the testing set.
In both validation exercises, we compared the performance of MINTS with the MICE PMM, pan Fixed Effects, pan Random Effects, Amelia TS, Amelia CS, and Amelia TSCS methods that were described in Section 4.4.

### v2

@@ 4.4 Model Implementation
For each replication of each experiment, we created 40 imputations using the MINTS method by running 10 chains of the MCMC algorithm. The bounds of the model for $\mathbf{X}$ were set as $X_{low}=0$ and $X_{up}=100$, and the bounds of the model for $\mathbf{Y}|\mathbf{X}$ were set as $Y_{low}=0$ and $Y_{up}=\text{min}(X_{c,t},60)$. During the estimation phase, the MCMC algorithm was run for enough iterations to ensure convergence of all imputation model parameters. The number of iterations differed across experiments, but ranged from 10,000 to 25,000 iterations per chain. During the imputation phase, an additional 4,000 iterations was run for each chain and four iterations from each chain were selected as the imputed values. The iterations selected as the final imputed values were chosen to be 1,000 iterations apart to ensure autocorrelation was close to zero following the procedure described in Section 3.3.2.
We compared the MINTS method to six models based on existing methods for multiple imputation. As the existing methods assume a linear relationship between variables with missing data, the simulated data was transformed prior to creating multiple imputations using a Box-Cox transformation. For each of the existing methods, $\mathbf{Y}$ was first transformed using the cube root as $\mathbf{Y}^{*}=\mathbf{Y}^{1/3}$. Imputations were created on the transformed scale, and the resulting multiply imputed values of $\mathbf{Y}^{*}$ were transformed back to the original scale of $\mathbf{Y}$.
@@
Three of the existing multiple imputation models considered are based on the MICE methodology as implemented in the R package “mice” (van Buuren and Groothuis-Oudshoorn (2011)).222mice version 3.14.0 used MICE uses the FCS algorithm, which allows for the specification of separate univariate conditional models for each variable with missing data. The univariate conditional models include all available variables as predictors, for example, the model for $\mathbf{X}$ includes $\mathbf{Y}$, year, and country as predictors. We first considered the default imputation method for continuous data in the mice package, which is predictive mean matching. We refer to this model as MICE PMM. Unlike the other approaches considered, MICE PMM does not explicitly account for the hierarchical structure of the data but is included as a baseline for comparison.
@@
We evaluated two models within the MICE framework that account for the hierarchical structure of the data using the method of Schafer (1997b) and Schafer and Yucel (2002), referred to as the pan method following its implementation in the R package “pan” (Zhao and Schafer (2023)). The function mice.impute.2l.pan within the mice package uses a Gibbs sampler to estimate the conditional linear mixed effects model with homogeneous within-group variances for each variable with missing data given the other variables. We evaluated a model that includes a country-specific intercept and fixed effects for all covariates in the imputation model. We refer to this model as pan Fixed Effects. For example, imputed values for $X_{c,t}|Y_{c,t}$ are modeled with a country-specific intercept, a fixed effect of year, a fixed effect of $Y_{c,t}$, and a homogeneous normally distributed error term. We also considered the pan Random Effects model, which adds random effects of all covariates to the model used in pan Fixed Effects.
@@
We evaluated three models using the Amelia II methodology as implemented in the R package “Amelia” (Honaker and King (2010)). [^3] Amelia version 1.80 used For complete data $(\mathbf{X},\mathbf{Y})$, Amelia assumes the joint distribution is multivariate normal. The parameters of the joint distribution are estimated using a combination of bootstrapping and an EM algorithm. Amelia is designed specifically for imputation of hierarchical time series data, which Honaker and King refer to as time series cross-sectional data, through modeling features such as smooth trends over time and allowing for country-specific effects. For our comparisons, we evaluated three implementations of Amelia that were chosen to use the simplest form of the time-series-cross-sectional modeling features in Amelia: a time-series (TS) model, a cross-sectional (CS) model, and a time-series-cross-sectional (TSCS) model. All three of the imputation models are constructed by adding terms to the default multivariate normal joint model for $(\mathbf{X},\mathbf{Y})$. In the Amelia TS model, a linear effect of time is added. In the Amelia CS model, a country-specific intercept term is added. In the Amelia TSCS model, a country-specific intercept term and a country-specific linear effect of time are added.
@@
We also considered a data augmentation approach that extends the MINTS method to enable estimation of the parameters of the analysis model through the decomposition
$p(\mathbf{Z},\mathbf{Y},\mathbf{X}|\bm{\theta}_{Z},\bm{\theta}_{Y},\bm{\theta}_{X})=$ $p(\mathbf{Z}|\mathbf{Y},\mathbf{X},\bm{\theta}_{Z})p(\mathbf{Y}|\mathbf{X},\bm{\theta}_{Y})p(\mathbf{X}|\bm{\theta}_{X}),$
where $\bm{\theta}_{Z}$ denotes the parameters of the analysis model. The data augmentation approach uses the same models for $\mathbf{Y}|\mathbf{X}$ and $\mathbf{X}$ as MINTS, while the model for $\mathbf{Z}|\mathbf{Y},\mathbf{X}$ is given by the analysis model used in the validation exercise. We abbreviate the data augmentation approach based on MINTS as the MINTS DA method. MINTS DA can be used to directly sample from the posterior distributions for the parameters of the analysis model rather than needing to rely on the multiple imputation framework. Medians of the posterior distributions for parameters of interest were used for point estimates, while the 0.025th and 0.975th quantiles of the posterior distributions were used for interval estimates. As the estimates of the parameters of interest are not constructed using Rubin’s pooling rules, we did not calculate estimates of the Fraction of Missing Information. We note that MINTS DA is considered only for comparison purposes as a method that directly incorporates information about $\mathbf{Z}$ when handling missing data in $\mathbf{Y}$ and $\mathbf{X}$. In practice, the data augmentation approach is not possible to implement in our motivating setting where the substantive analysis of interest and imputation of missing data are conducted independently, and information about $\mathbf{Z}$ and $\mathbf{X}$ are not shared between analyst and imputer.
@@ 4.5 Analysis Model Validation Results
For the linear regression analysis model validation, we evaluated how well each multiple imputation method performs for estimation of $Q=\omega_{1}$, the regression coefficient on $\mathbf{Y}$ in the linear regression of $\mathbf{Z}$ on $\mathbf{Y}$. Table 1 summarizes the results of this validation exercise.
Overall, we found MINTS results in the best balance between MAE, coverage, and FMI. Out of the multiple imputation methods considered, MINTS has the smallest MAE in all experiments except MAR 10% and MAR 40%, where MINTS has the second smallest MAE behind MICE PMM. MINTS has reasonably close to nominal coverage for 95% intervals at the 10% and 40% rates of simulated missingness, but suffers from undercoverage at the 80% rate. The Amelia TSCS method has closer to nominal coverage than MINTS in the MCAR 80% and MNAR 80% experiments at the expense of higher MAE, but MINTS has the closest to nominal coverage in the MAR 80% experiment.
MINTS has the smallest FMI in eight out of nine experiments. As expected, FMI is much larger for all methods at the 80% rate, with FMI $>50\%$ for several methods. This is higher than is typical for the setting of sample survey data that multiple imputation was originally designed for, where FMI is usually $\leq 30\%$ (Rubin 2007). Given the high FMI, it is thus unsurprising how poorly all multiple imputation methods perform in the experiments at the 80% rate. In particular, the spline estimation used in MINTS suffers from overfitting at the 80% rate due to the small number of complete cases available to estimate the splines $f$ and $h$, with some replications having less than 15 complete cases.
None of the existing multiple imputation methods perform consistently well across experiments for estimation of $\omega_{1}$. While MICE PMM can outperform the methods explicitly designed for hierarchical time series data at the lower rates of missingness, the performance of MICE PMM suffers greatly at the highest rate of missingness. The pan and Amelia methods perform similarly to one another at the 10% rate, but all have larger MAE than MICE PMM. Performance of the pan and Amelia methods generally is best under MCAR, with pan Random Effects and Amelia TSCS performing the best of the group in terms of MAE and coverage.
Multiple imputation using the MINTS method generally has better performance than the data augmentation approach based on MINTS at the lower rates of simulated missingness. However, the data augmentation approach has smaller MAE than all multiple imputation methods at the highest rate of simulated missingness. The multiple imputation methods all have substantial undercoverage in the MAR 80% and MNAR 80% experiments, while the data augmentation approach maintains close to nominal coverage. These findings suggest that omission of $\mathbf{Z}$ from the imputation model is likely a major source of bias and undercoverage for the multiple imputation methods when there is a small number of observed values.
@@
Simulated Method MCAR MAR MNAR Missingness Rate MAE Cvg FMI MAE Cvg FMI MAE Cvg FMI 10% MICE PMM 0.85 100.0 5.1 0.38 100.0 1.6 0.49 100.0 2.6 pan Fixed 1.36 100.0 8.6 1.69 100.0 1.9 1.06 100.0 3.2 pan Random 1.08 100.0 7.6 1.28 100.0 1.4 0.77 100.0 2.3 Amelia TS 1.60 100.0 10.9 1.89 100.0 2.6 1.13 100.0 3.8 Amelia CS 1.38 100.0 7.9 2.26 100.0 4.4 1.29 100.0 3.7 Amelia TSCS 1.28 100.0 8.1 1.40 100.0 1.3 0.70 100.0 2.1 MINTS 0.29 100.0 0.5 0.56 100.0 0.5 0.29 100.0 0.4 MINTS DA 1.00 100.0 0.67 100.0 0.97 100.0 40% MICE PMM 12.60 40.0 36.6 2.22 100.0 13.9 8.45 77.8 28.4 pan Fixed 6.87 91.2 33.2 10.13 15.3 14.9 8.57 66.7 22.6 pan Random 5.33 94.3 36.9 6.54 88.8 12.1 4.19 99.9 15.9 Amelia TS 7.18 94.3 29.0 10.63 13.9 17.1 7.83 82.4 19.9 Amelia CS 14.25 35.2 26.8 12.11 33.4 19.0 12.17 21.9 21.1 Amelia TSCS 4.70 93.6 38.3 6.96 81.2 17.1 3.50 99.4 17.2 MINTS 1.21 100.0 3.9 3.02 99.9 4.4 1.49 100.0 2.8 MINTS DA 2.38 91.3 1.80 98.7 2.21 98.5 80% MICE PMM 58.07 0.1 60.1 24.73 14.1 52.7 56.68 0.0 53.9 pan Fixed 32.27 8.6 65.2 29.84 0.0 47.7 37.56 0.0 58.8 pan Random 42.19 34.6 87.5 38.34 17.5 84.3 45.67 47.5 86.1 Amelia TS 20.38 69.3 54.4 33.08 0.0 42.3 24.87 3.7 39.5 Amelia CS 91.98 0.0 72.7 70.19 0.0 79.0 63.14 0.0 64.5 Amelia TSCS 30.58 91.0 93.8 40.11 25.2 89.1 27.10 75.4 89.1 MINTS 7.62 81.0 38.7 13.14 47.5 47.0 9.71 62.6 35.6 MINTS DA 4.05 88.8 4.53 99.4 3.36 96.9
We also conducted a simulation study for data simulated to follow a linear relationship. For the linear simulated data, we considered two uncongenial and two congenial analysis models. Although our primary focus is on the uncongenial setting, the analysis model validation using congenial analysis models and linear simulated data allows us to compare the performance of MINTS with the existing imputation methods in an “ideal” setting for multiple imputation. We found that the existing imputation methods perform better in the analysis model validation for the linear simulated data compared to the nonlinear simulated data. However, for estimation of uncongenial analysis models, we found MINTS still outperforms the existing methods in the linear setting at the 10% and 40% rates of simulated missingness. All imputation methods were found to perform well for estimation of congenial analysis models using linear simulated data, with no method consistently standing out from the others across experiments. Details of the simulation study using linear simulated data can be found in Sections 5 and 6 of the Supplementary Material (Liu and Raftery (2025)).
@@ 5 Application to Enrollment Data
We further evaluated the performance of MINTS by revisiting the secondary school enrollment data that was described in Section 2. We conducted two validation exercises using the school enrollment data by simulating additional country-years as missing. In the first validation exercise, we evaluated how well MINTS performs for estimation of parameters of interest $Q$ for uncongenial analysis models. This is analogous to the analysis model validation that was conducted for the simulation study.
In the second validation exercise, we evaluated the predictive performance of MINTS for predicting left-out observations of NER. Out-of-sample validation for prediction of individual missing values is less frequently conducted for multiple imputation methods compared to analysis model validation, but has been considered by Gelman, King and Liu (1998); Honaker and King (2010); Nguyen, Carlin and Lee (2017); among others. Although the primary goal of multiple imputation is to create valid estimates of parameters of interest in the presence of missing data rather than recovering the missing values (Rubin (1996)), the prediction of individual missing values can still be of great interest in practice. A multiple imputation method that can perform well for prediction of individual missing values and for creating multiply imputed estimates of quantities of interest is thus of increased utility. We refer to this second validation exercise as “out-of-sample validation.”
Finally, we applied the MINTS method to the full school enrollment data set without simulating any additional missing values and created 40 multiple imputations for the country-years of NER that are missing in the original data set. We also estimated a substantive analysis model of interest using the multiply imputed values of NER created using MINTS and compared these results with the estimates obtained using the other methods considered in the validation exercises.
@@ 5.1 Validation Procedure
For both validation exercises using the school enrollment data, we considered eight experiments. Parameters varied in both validation exercises were the rate of simulated missingness (10%, 40%, 80%) and the missing data mechanism (MCAR, MAR, and MNAR), where the same missing data mechanisms used in the simulation study and described in Section 3 of the Supplementary Material (Liu and Raftery (2025)) were also used for the enrollment data. We did not consider the MNAR 80% experiment for the enrollment data as this resulted in the majority of countries having no observations for NER or GER.
@@
For the analysis model validation, each experiment was replicated $N_{rep}=100$ times following an analogous procedure as was described in Section 4.3 for the simulation study. A major difference in the analysis model validation procedure for the enrollment data is the need to distinguish between the country-years that start out as missing and the country-years that are simulated as missing, where only the country-years that are simulated as missing are used for evaluation purposes. To facilitate this distinction, in each replication of each experiment we separated the enrollment data into “started-as-missing”, “observed”, and “simulated-as-missing” sets. The observed set is the training set for model estimation, while the simulated-as-missing set is the testing set for validation. Imputed values were still created for the started-as-missing set, but we are unable to evaluate the performance of these imputed values as the true values for the started-as-missing set are unknown.
For the out-of-sample validation exercise, we considered one replication of each experiment. We used the same distinction between the country-years that started as missing and the country-years that were simulated as missing, where only the country-years that were simulated as missing are included in the testing set for validation. The out-of-sample validation exercise compares performance metrics for prediction of missing values for NER averaged over all country-years in the testing set.
In both validation exercises, we compared the performance of MINTS with the MICE PMM, pan Fixed Effects, pan Random Effects, Amelia TS, Amelia CS, and Amelia TSCS models that were described in Section 4.4. For the analysis model validation exercise, we additionally compared the multiple imputation methods with the data augmentation approach described in Section 4.4.